OpenStudy (anonymous):

What is the area of a regular hexagon with a side length of 9 inches and an apothem length of 7.8 inches? A. 175.8 square inches B. 186.5 square inches C. 210.6 square inches D. 225.4 square inches

OpenStudy (anonymous):

@iGreen

OpenStudy (anonymous):

6*0.5*9*7.8 = 210.6 in^2 its C

OpenStudy (igreen):

@MAli13chineta Don't give out answers.

OpenStudy (anonymous):

I just keep forgetting the formula or i get it confused with others

OpenStudy (igreen):

Formula for area of a hexagon is \(\dfrac{5}{2}sa\), where s = side, and a = apothem. \(\dfrac{5}{2}(9)(7.8)\) Multiply: \((22.5)(7.8)\) Can you multiply those two?

OpenStudy (igreen):

If you forget the formula just Google it :)

OpenStudy (anonymous):

very true having a blonde moment geometry is weakest subject

OpenStudy (anonymous):

is my

OpenStudy (igreen):

Lol

OpenStudy (igreen):

Can you multiply 22.5 and 7.8? @raeluvpop